A firm operating in a perfectly competitive market has to sell

all its output at the price of $10 per unit. Its marginal cost

function is given by Q + 4 and the total fixed cost is 1.

Determine;

A. The profit maximizing output level.

B. The level of supernormal profit if any.


1
Expert's answer
2021-06-03T13:11:57-0400

A. In a perfect competition market;

Max profit is attained at a level of Output (Q) where MR = MC = P = D i.e. Marginal revenue equals to Marginal Cost, Price or Demand.

But MC = Q + 4 and P = 10

10 = Q + 4

Hence Q = 10 - 4 = 6

"\therefore" The profit maximizing output level is 6 units.


B. In a perfect competition market; firms are price takers, and therefore maximum profits can only be attained at the level of ouput where MR=MC = P beyond which costs will exceed revenues and firms will start making losses.Any additional units sold, revenues will increase by the same amount equal to the market price.

For instance, beyond 6 units let's say 7 units, the MR = $10 while MR = Q + 4 = 7 + 4 = $11; TR = 7 "\times" 10 = $70 and TC = 1 + (11"\times"7) = 78; Loss will be 78 - 70 = $8

So, super-normal profits are not attainable in a perfect competition market.